Maryam Inc. produces leather bound books. The market is highly competitive, with books currently selling for R600. Maryam’s total and marginal cost curves are: TC = 3 000 + 100Q2 MC = 200Q Calculate Maryam Inc.’s profit maximising quantity. Is the firm earning a profit? / Mary


1
Expert's answer
2021-10-24T18:15:45-0400

"TR=600Q"

"MR=600"

"MC=200Q"

To find profit maximizing quantity:

"MR=MC"

"600=200Q"

"Q=3"

Profit maximizing quantity=3.


"TC=3000+100Q^2"

"TC=3000+100(3^2)=3900"

"TR=600Q=600\times3=1800"

"Profit\/loss=TR-TC"

"=1800-3900=-2100"

Thus, the firm is making a loss.